# Echoes of $R^{3}$ modification and Goldstone preheating in the CMB-BAO landscape

### v1

@@ Abstract
Discussion and Summary.— We showed that $R^{3}$ can make the single-field-like regime of the $R^{2}$-Higgs model compatible with observations. The parameter space that accounts for the observed high $n_{s}$ also induces strong Goldstone preheating, which in turn helps match the reference CMB scale to inflationary dynamics. Using two representative parameter sets, we find that in the $R^{2}$-like regime, with nonminimal coupling $\xi_{H}\gtrsim 1.5$, a $\xi_{c}\sim 10^{-14}$ is sufficient to trigger preheating. While a nonvanishing $\xi_{c}$ can still account for the high $n_{s}$, Goldstone and Higgs preheating require $\xi_{H}\gtrsim 1.5$. We primarily focused on the $R^{2}$-like regime; however, in the Higgs-like regime with a larger $\xi_{H}$, preheating is faster and matching becomes easier. This novel connection between the CMB-BAO tension, higher-order curvature effects, and preheating in $R^{2}$-Higgs inflation has not been addressed in the literature. We also note that our study leaves room for improvement. Primary uncertainties arise from unaccounted condensate and perturbation decays, which may require going beyond the linear-order approximation and including non-equilibrium effects.
@@
Appendix.— The mass matrix $\mathcal{M}^{I}_{\ L}$ and $m_{\mathrm{eff},(I)}^{2}$ are
$\mbox{and}~m_{\mathrm{eff},(I)}^{2}(\tau)=\mathcal{M}^{I}_{~~I}-\frac{1}{6}R_{E}G^{I}_{~I},$ (14)
where $R_{E}=-(2H^{2}+\dot{H})$. The $\mathcal{K}_{Z}$ and $\Upsilon$ in Eq. (7) are
The required BD initial condition to solve Eq. (7) is given as: $s_{k}=e^{-ik\tau}/(\sqrt{2k\Delta_{(\phi_{2})}})$ and $s_{k}^{\prime}$. The $\Delta_{(\phi_{2})}$ factor in the denominator appears due to the presence of the second friction term in Eq. (7).

### v2

@@ IV Inflationary observables and scale matching
V Discussion and Summary
We showed that $R^{3}$ can make the single-field-like regime of the $R^{2}$-Higgs model compatible with observations. The parameter space that accounts for the observed high $n_{s}$ also induces strong Goldstone preheating, which in turn helps match the reference CMB scale to inflationary dynamics. Using two representative parameter sets, we find that in the $R^{2}$-like regime, with nonminimal coupling $\xi_{H}\gtrsim 1.5$, a $\xi_{c}\sim 10^{-14}$ is sufficient to trigger preheating. While a nonvanishing $\xi_{c}$ can still account for the high $n_{s}$, Goldstone and Higgs preheating require $\xi_{H}\gtrsim 1.5$. We primarily focused on the $R^{2}$-like regime; however, in the Higgs-like regime with a larger $\xi_{H}$, preheating is faster and matching becomes easier. It should also be noted that, unlike pure Higgs inflation, here the unitarity cut-off is restored up to the Planck scale, see e.g. Ref. [57] and, the produced Goldstone bosons do not violate the unitarity. This novel connection between the CMB-BAO tension, higher-order curvature effects, and preheating in $R^{2}$-Higgs inflation has not been addressed in the literature. We also note that our study leaves room for improvement. Primary uncertainties arise from unaccounted condensate and perturbation decays, which may require going beyond the linear-order approximation and including non-equilibrium effects.